OpenStudy (anonymous):

what is the derivative of the integral from 2 to 1/X of sin^4 t dt

OpenStudy (anonymous):

\[f(x)=\int_2^x\sin^4(t)dt\] \[f'(x)=\sin^4(x)\] because the derivative of the integral is the intergrand thefore \[f(\frac{1}{x})=\int_0^{\frac{1}{x}}\sin^4(t)dt\] \[f'(\frac{1}{x})=\frac{-1}{x^2}\sin^4(\frac{1}{x})\] by the chain rule

OpenStudy (anonymous):

typo as i changed the lower limit of integration but it does not change the answer

OpenStudy (anonymous):

good morning all ( at least morning here)

OpenStudy (anonymous):

Use Leibniz integral rule http://en.wikipedia.org/wiki/Leibniz_integral_rule to get your answer and Good evening friends (here)

OpenStudy (anonymous):

@ellex11 you got this?

OpenStudy (anonymous):

@satellite73 I think you lost the OP when you posted the answer without engaging any discussion what-so-ever. I don't believe this is a race to see how fast WE can answer the question, rather how WELL we can help the OP. just a thought.

OpenStudy (anonymous):

It's actually very easy applying Leibnitz's theorem, \[F(x) = \large\int \limits_2 ^ {\frac{1}{x}} \sin ^4 t \; dt \] Applying Leibnitz's theorem, we get, \[ F'(x) = \sin^4 \left(\frac{1}{x} \right) \cdot \frac d{dx} \frac{1}{x} - \sin^4 \left(2 \right) \cdot \frac d{dx} (2) \] \[ = -\frac 1 {x^2} \sin^4 \left(\frac{1}{x} \right)\]

OpenStudy (anonymous):

true, but kind of weird strange since you have a constant

OpenStudy (anonymous):

Sat, for a monovariant function \(g\): \[ {d\over dx}\, \int_{f_1(x)}^{f_2(x)} g(t) \,dt = g(f_2(x)) {f_2'(x)} - g(f_1(x)) {f_1'(x)} \] can't we assume \( f_1(x) = 2 \)?

OpenStudy (anonymous):

yes yes, of course it works.

OpenStudy (anonymous):

Aha!, thanks :)